OpenStudy (anonymous):

Solve -11(8 – x) = x – 10(5 – x). Please show all of your work

OpenStudy (gw2011):

This problem has no solution because of the following: -11(8-x)=x-10(5-x) -88+11x=x-50+10x -88+11x=11x-50 11x-11x=88-50 0 does not equal 38. Therefore, there is no solution to this problem as it is stated.